OpenStudy (anonymous):

The function f(t) = 4t^2 - 8t + 8 shows the height from the ground f(t), in meters, of a roller coaster car at different times t. Write f(t) in the vertex form a(x - h)^2 + k, where a, h, and k are integers and interpret the vertex of f(t).

OpenStudy (anonymous):

@Harsha19111999

OpenStudy (harsha19111999):

\[y-4=4(t-1)^2\]

OpenStudy (harsha19111999):

\[y=4(t-1)^2+4\]

OpenStudy (anonymous):

f(t) = 4(t - 1)2 + 4; the minimum height of the roller coaster is 1 meter from the ground f(t) = 4(t - 1)2 + 4; the minimum height of the roller coaster is 4 meters from the ground

OpenStudy (anonymous):

which one @Harsha19111999

OpenStudy (harsha19111999):

It must be minimum and maximum

OpenStudy (harsha19111999):

Cause you have 2 cases